OpenStudy (anonymous):

How many percent faster is 15 min than 19 min? And how many percent slower is 19 min than 15 min?

OpenStudy (baldymcgee6):

a percent is out of 100, so if you divide 15/19 * 100 it will give you a percent

OpenStudy (anonymous):

A percent of what? I don't get it.

OpenStudy (baldymcgee6):

15/19=0.789 0.789*100=78.9%

OpenStudy (anonymous):

And what does it answer? Sorry... I'm slow, I know.

OpenStudy (baldymcgee6):

15mins is 78.9% faster than 19mins

OpenStudy (anonymous):

That doesn't sound right. Sorry. But it doesn't. It's too much, isn't it? Since 50 % of 19 is 9,5, and that proves that it isn't right. If you get how I mean.

OpenStudy (anonymous):

Or something.

OpenStudy (baldymcgee6):

and 78.9% of 19 is 15

OpenStudy (anonymous):

Now I get it. Thank you.

OpenStudy (anonymous):

But.. How do I solve the first question then?

OpenStudy (baldymcgee6):

okay so, lets get this straight. 15/19 * 100 = 78.9% right? agree?

OpenStudy (baldymcgee6):

and 19/15 *100 = 126.7%

OpenStudy (baldymcgee6):

so 15min is 126.7% faster than 19min and 19 min is 78.9% slower than 15 min.

OpenStudy (anonymous):

You're smart. O.o

OpenStudy (baldymcgee6):

does it make sense?

OpenStudy (anonymous):

Yesh.